OpenStudy (anonymous):

I am getting the incorrect answer for determining the eigenvectors of a matrix. I will post my work in the next post.

OpenStudy (amistre64):

eugene vectors are such that: Ax = Ix, or succinctly; (A-I)x = 0 and im sposing we start with the characteristic equation of A-I ... if memory serves

OpenStudy (amistre64):

curious on what your matrix is tho :)

OpenStudy (anonymous):

The matrix in question is \[\left[\begin{matrix}1 & 2 \\ 2 & 3\end{matrix}\right]\] It's easy to work out the eigenvalues with the formula \[\left| A - \lambda I \right| = 0\] They are \[2+\sqrt{5}\] and \[2-\sqrt{5}\] However, when I try to determine the eigenvectors, I get the wrong answer. Please note that the eigenvectors are unit vectors, so we can use \[v _{1}^{2} + v_{2}^{2} = 1\] The correct eigenvectors are \[\left[\begin{matrix}.5257 & .8507 \\ -.8506 & .5257\end{matrix}\right]\] Attached is my work which gives me the incorrect answer. I would like to understand what I'm doing wrong.

OpenStudy (beginnersmind):

Where do the values \(2+\sqrt{5} \text{ and } 2-\sqrt{5}\) come from?

OpenStudy (beginnersmind):

 Sorry, too lazy to use LaTex. Solving the last line should give the value of lambda. Is this what you did?